OpenStudy (anonymous):

Michelle uses a credit card with a 23.2% APR compounded monthly to pay for car repairs totaling $790.59. She can pay $330 per month on the card. What will the total cost of this purchase be? $974.01 $790.59 $818.07 $948.66

OpenStudy (anonymous):

I found a formula that might help you... A=P(1+r/n)^nt P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year oh by the way I think this question should be moved to a different section. Perhaps mathematics.

OpenStudy (amistre64):

trying to find "t" eh...\[B_n=B_ok^{12*t}-P(\frac{1-k^{12*t}}{1-k})\] \[0=B_ok^{12*t}-P(\frac{1-k^{12*t}}{1-k})\] \[B_ok^{12*t}=P(\frac{1-k^{12*t}}{1-k})\] \[\frac{B_o(1-k)}{P}k^{12*t}=1-k^{12*t}\] \[\frac{B_o(1-k)}{P}=k^{-12*t}-1\] \[log\left(\frac{B_o(1-k)+P}{P}\right)=-12*t~log(k)\] \[\frac{-1}{12log(k)}log\left(\frac{B_o(1-k)+P}{P}\right)=t\] maybe

OpenStudy (amistre64):

bah .... need to see if i missed the algebra along the way

OpenStudy (amistre64):

the wolfs being a pain .....

OpenStudy (amistre64):

when i run it with an excel setup; we get 2 payments + 158.07 to finish it 790.59 +330.00 +330.00 +158.07 -------- 1608.66

OpenStudy (amistre64):

lol, and when i dont charge double ... we get 818

OpenStudy (amistre64):

790.59(1.0193) - 330.00 = 475.87 475.87(1.0193) - 330.00 = 155.07 155.07(1.0193) - 158.07 = 0